OpenStudy (anonymous):

Find the value of x and y in the following equation. 66-2x=5y 7y+18=-4x

OpenStudy (anonymous):

-2x-5y=-66 4x+7y=-18 Multiply the top equation by 2: -4x-10y=-132 4x+7y=-18 Add the equations together: -3y=-150 y = 50 Substitute y=50 back into original (I'll just use the first equation): -2x-5(50)=-66 -2x-250=-66 -2x=184 x = -92 Therefore: x= -92 y= 50

OpenStudy (anonymous):

Thanks :)